# Excess cataloged X-ray and radio sources at galaxy-cluster virial shocks

### 26 Feb 2024 v1

@@ Abstract
We detect a highly significant excess of X-ray (2RXS) and radio (NVSS, GMRT, VLSSr) catalog sources when stacked around MCXC galaxy clusters and groups, narrowly confined within $\lesssim 100\mathrm{\,kpc}$ of the $\sim 2.4R_{500}$ virial shock radius (inferred from continuum stacking), with similar X-ray ($\sim 4\sigma$ for $443$ clusters) and radio ($\sim 4\sigma$ for $485$ clusters) characteristics ($>5\sigma$ joint). The excess sources show 10101010100100100100 kpc scales, $L_{X}(0.1\mbox{--}2.4\mbox{ keV})\simeq 10^{42\mbox{\scriptsize--}43}\mathrm{\,erg\,s^{-1}}$ or $\nu L_{\nu}(\nu=1.4\mathrm{\,GHz})\simeq 10^{40{\mbox{\scriptsize--}}41}\mathrm{\,erg\,s^{-1}}$ luminosities, and a preferentially radial radio-polarization. The narrow localization and properties of the excess identify these sources not as AGN, often invoked speculatively for excess X-ray sources at cluster outskirts, but rather as infalling gaseous clumps interacting with the virial shock, probably galactic halos and possibly outflow remnants. The local excess of such discrete, radio-to-$\gamma$-ray sources around an object can probe its virial shock also at high redshifts and sub-cluster scales.
@@ 1 Introduction
More recently, a $\sim 3\sigma$ excess of Chandra X-ray sources was reported [koulouridis2019high] in a fairly narrow, $2.52<\tau<$ 2.5 range of normalized radii around a small sample of five $z\sim 1$, mostly disturbed clusters. This signal is dominated by two clusters, each with about ten excess $2.52.0<\tau<$ 2.5 sources with respect to neighbouring bins. The excess sources show $10^{42.5}{\rm{\,erg\,s^{-1}}}\lesssim L_{0.5\mbox{ \scriptsize keV}}^{8.0\mbox{ \scriptsize keV}}\lesssim 10^{44}{\rm{\,erg\,s^{-1}}}$ luminosities, where energy or frequency subscripts (superscripts) denote low (high) band limits, henceforth. These sources were again attributed to enhanced AGN activity, due to excessive galaxy mergers or gas stripping.
@@
Therefore, a sharp spike in the density of sources would indicate a localized, sudden change in the environment. Interestingly, the structure formation, or virial, shock of a galaxy cluster, long anticipated to lie in the $32<\tau<$ 3 range according to cosmological simulations [e.g., KeshetEtAl03, SchaalSpringel15], provides a natural candidate for such a sharp environmental transition. Indeed, as an infalling object crosses the virial shock (VS), it should see the ambient pressure suddenly jump by orders of magnitude as the infalling surrounding plasma is violently slowed down and heated, accompanied by the generation of magnetic fields and shock-accelerated relativistic particles (cosmic rays, CRs) carrying a substantial fraction of the elevated thermal energy. While crossing a VS is unlikely to trigger an AGN, and the processes by which it could lead to the emergence of bright excess X-ray or radio sources may be a-priori unclear, the VS could explain at least the position and sharpness of the signal.
@@
The Supplementary material provides supporting and auxiliary information regarding the spectra of the excess X-ray sources (§A), limitations on source size estimates (§B), the evaluation of excess confidence levels (CLs) under Poisson statistics (§C), various convergence and sensitivity tests (§D), additional properties of the excess sources (§E), and the absence of a significant VS signal from 2RXS–NVSS pairs (§F) or identified AGN (§G).
@@ 2.1 Cluster catalog and selection criteria
    Additional cuts are imposed on the Galactic latitude $|b|$, to avoid the bright Galactic plane, and on the declination Dec. These cuts are specific to each source catalog, so are detailed below.
@@ 2.2 X-ray source catalog and cuts
Figure 1: Hammer-Aitoff all-sky projection of source density (cubehelix color map [Green11_Cubehelix]) and cluster sample (orange circles of radius $5{\theta_{500}}$), for X-ray (left panel; 2RXS) and radio (right panel, NVSS; note diminished exposure at low declinations with no $\mbox{Dec}<-40^{\circ}$ sources) analyses. Also shown are the cuts of the Galactic plane (dotted yellow lines), the Fermi bubbles (dashed red curves), and Loop-I (dot-dashed cyan).
@@
[…]
@@
[…]
@@
  • ${}^{a}$ — Energy flux (specific energy flux) in X-rays (radio). The cuts are explained in the text.
    ${}^{b}$ — After cuts.
We define the X-ray energy flux $F_{X}\equiv F_{0.1\mbox{ \scriptsize keV}}^{2.4\mbox{ \scriptsize keV}}$ as the specific flux $F_{\epsilon}$ integrated over the ROSAT band. The catalog provides flux estimates based on three different emission models: a power-law spectrum, the optically-thin thermal plasma emission model Mekal [MeweEtAl86], and optically thick, black-body emission. In most or all relevant cases, the model fits are not sufficiently accurate to identify which emission model is most appropriate. Therefore, we henceforth adopt the power-law estimate of $F_{X}$, which is a simpler fit available for the largest number of sources; our results are not sensitive to this choice (see §D). We nevertheless consider the fitted source temperatures, as well as the fitted photon power-law index $\Gamma_{X}$ defined by $F_{\epsilon}\propto\epsilon^{1+\Gamma_{X}}$.
Figure 2 (left panel) shows the energy flux histogram of X-ray sources in the analyzed sky region. The 2RXS catalog is dominated by foreground sources brighter than the stacked clusters themselves. To reduce foreground confusion, we impose an upper flux limit that removes sources whose luminosities would exceed the typical, $L_{X}\simeq 10^{45}{\rm{\,erg\,s^{-1}}}$ luminosity of the entire ICM even around the most distant ($z_{\rm max}\simeq 0.23$) clusters in our sample, thus retaining only $F_{X}<F_{X}^{*}\equiv 7\times 10^{-12}{\rm{\,erg\,s^{-1}\,cm^{-2}}}$ sources. Note that this $F_{X}^{*}$ cutoff is already quite close to the 2RXS $\sim 10^{-13}{\rm{\,erg\,s^{-1}\,cm^{-2}}}$ sensitivity threshold [2RXS_paper]. In §D, we show that our results are not sensitive to the precise choice of flux cut; for example, a high $F_{X}^{*}=10^{-10}{\rm{\,erg\,s^{-1}\,cm^{-2}}}$ changes the significance of excess virial sources by only $\sim 2\%$.
@@ 2.3 Radio source catalog and cuts
The NVSS catalog covers the northern sky at declinations $\mbox{Dec}>-40^{\circ}$ (J2000). Figure 1 shows an apparent decline in $n_{\nu}$ with decreasing Dec even somewhat north of this declination. Therefore, we impose a corresponding cut on the declination of clusters used for radio stacking, keeping only those with $\mbox{Dec}>-35^{\circ}$.
@@
A histogram of the specific flux of NVSS sources is shown in the right panel of figure 2. This catalog is dominated by faint background sources (in contrast to the bright foreground source of 2RXS), due in part to star-forming galaxies, whose luminosity function peaks near $\nu L_{\nu}\sim 10^{39}{\rm{\,erg\,s^{-1}}}$ [e.g., radio_Lum]. We therefore impose a lower specific-flux limit that removes sources with luminosities lower than this peak luminosity even around the nearest ($z_{\rm min}\simeq 0.017$) clusters in our sample, thus retaining only sources of $F_{\nu}\gtrsim F_{\nu}^{*}\equiv 100\mbox{ mJy}$. In §D, we show that changing the radio flux cutoff by $50\%$ has only a $\sim 15\%$ effect on the significance of our results, and our choice is conservative.
@@ 3 Stacking analysis method
$n(\tau,{\mathsf{c}})=\frac{\mathcal{N}(\tau,{\mathsf{c}})}{\Delta\omega(\tau,{\mathsf{c}})}\mbox{ ,}$ (1)
@@
$\Delta\omega(\tau,{\mathsf{c}})\equiv\frac{\Delta\Omega(\tau,{\mathsf{c}})}{\theta_{500,{\mathsf{c}}}^{2}}$ (2)
@@
$f(\tau,{\mathsf{c}})=\frac{\mathcal{F}(\tau,{\mathsf{c}})}{\Delta\omega(\tau,{\mathsf{c}})}\mbox{ ,}$ (3)
@@
$\delta(\tau,{\mathsf{c}})\equiv\frac{\mathcal{N}(\tau,{\mathsf{c}})-\mathcal{F}(\tau,{\mathsf{c}})}{\mathcal{F}(\tau,{\mathsf{c}})}=\frac{n(\tau,{\mathsf{c}})-f(\tau,{\mathsf{c}})}{f(\tau,{\mathsf{c}})}\mbox{ .}$ (4)
@@
$S\simeq\frac{\mathcal{N}-\mathcal{F}}{\sqrt{\mathcal{F}\,}}\mbox{ .}$ (5)
@@
$S_{\mbox{\tiny{joint}}}\simeq\frac{1}{\sqrt{N_{\mbox{\tiny{cat}}}}}\sum_{i}^{N_{\mbox{\tiny{cat}}}}S_{i},$ (6)
@@ 3.1 Source-weighted co-addition
$\mathcal{N}(\tau)\equiv\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\mathcal{N}(\tau,{\mathsf{c}})\mbox{ ,}$ (7)
@@
$\Omega(\tau)\equiv\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\Delta\Omega(\tau,{\mathsf{c}})\mbox{ ,}$ (8)
@@
$n_{\rm\mbox{\tiny{SW}}}(\tau)=\frac{\mathcal{N}(\tau)}{\bar{\omega}(\tau)}\mbox{ ,}$ (9)
@@
$\sigma[n_{\rm\mbox{\tiny{SW}}}(\tau)]\simeq\frac{\sqrt{\mathcal{N}(\tau)}}{\bar{\omega}(\tau)}\mbox{ .}$ (10)
Source-weighted estimates such as $n_{\rm\mbox{\tiny{SW}}}(\tau)$ are somewhat skewed toward extended clusters or clusters in strong-field regions, which may harbour a larger number of sources and so may dominate the numerator of Eq. (9). This bias is offset in part by the comparable extents of all clusters in our sample, and by avoiding regions with known contaminations.
The above approach is applied analogously to field sources, modelled with some distribution $\mathcal{F}(\tau,{\mathsf{c}})$ based on the field region. In analogy with Eqs. (7)–(9), the co-added number of field sources is
$\mathcal{F}(\tau)\equiv\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\mathcal{F}(\tau,{\mathsf{c}})\mbox{ ,}$ (11)
@@
$f_{\rm\mbox{\tiny{SW}}}(\tau)=\frac{\mathcal{F}(\tau)}{\bar{\omega}(\tau)}\mbox{ .}$ (12)
The significance of a local excess of sources with respect to the field-only null hypothesis can now be estimated as in Eq. (5),
$S_{\rm\mbox{\tiny{SW}}}(\tau)=\frac{\mathcal{N}(\tau)-\mathcal{F}(\tau)}{\sqrt{\mathcal{F}(\tau)}}\mbox{ ,}$ (13)
@@
$\delta_{\rm\mbox{\tiny{SW}}}(\tau)=\frac{n_{\rm\mbox{\tiny{SW}}}(\tau)}{f_{\rm\mbox{\tiny{SW}}}(\tau)}-1\mbox{ ,}$ (14)
@@ 3.2 Cluster-weighted co-addition
$S(\tau,{\mathsf{c}})\simeq\frac{\mathcal{N}(\tau,{\mathsf{c}})-\mathcal{F}(\tau,{\mathsf{c}})}{\sqrt{\mathcal{F}(\tau,{\mathsf{c}})}}\mbox{ .}$ (16)
@@
$S_{\rm\mbox{\tiny{CW}}}(\tau)=\frac{1}{\sqrt{{N}_{c}}}\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}S(\tau,{\mathsf{c}})\mbox{ .}$ (17)
@@
$n_{\rm\mbox{\tiny{CW}}}(\tau)=\frac{1}{{N}_{c}}\sum_{{\mathsf{c}}=1}^{{N}_{c}}n(\tau,{\mathsf{c}})\mbox{ ,}$ (18)
@@
$\sigma[n_{\rm\mbox{\tiny{CW}}}(\tau)]=\frac{1}{{N}_{c}}\sqrt{\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\frac{n(\tau,{\mathsf{c}})}{\Delta\omega(\tau,{\mathsf{c}})}}\mbox{ .}$ (19)
@@
$f_{\rm\mbox{\tiny{CW}}}(\tau)=\frac{1}{{N}_{c}}\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}f(\tau,{\mathsf{c}})\mbox{ .}$ (20)
@@
$\delta(\tau,{\mathsf{c}})=\frac{n(\tau,{\mathsf{c}})}{f(\tau,{\mathsf{c}})}-1\mbox{ ,}$ (21)
@@
$\delta_{\rm\mbox{\tiny{CW}}}(\tau)=\frac{1}{{N}_{c}}\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\delta\left(\tau,{\mathsf{c}}\right)\mbox{ ,}$ (22)
@@
$\sigma\left[\delta_{\rm\mbox{\tiny{CW}}}(\tau)\right]=\frac{1}{{N}_{c}}\sqrt{\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\frac{n^{2}(\tau,{\mathsf{c}})}{f^{2}(\tau,{\mathsf{c}})\mathcal{N}(\tau,{\mathsf{c}})}}\mbox{ ,}$ (23)

### 19 Sep 2024 v2

@@ Abstract
We detect a highly significant excess of X-ray (2RXS) and radio (NVSS, GMRT, VLSSr) catalog sources when stacked around MCXC galaxy clusters and groups, narrowly confined within $\lesssim 100\mathrm{\,kpc}$ of the $\sim 2.4R_{500}$ virial shock radius (inferred from previous continuum stacking), with similar X-ray ($\sim 4\sigma$ for $443$ clusters) and radio ($\sim 4\sigma$ for $485$ clusters) characteristics ($>5\sigma$ joint). The excess sources show 10101010100100100100 kpc scales, $L_{X}(0.1\mbox{--}2.4\mbox{ keV})\simeq 10^{42\mbox{\scriptsize--}43}\mathrm{\,erg\,s^{-1}}$ or $\nu L_{\nu}(\nu=1.4\mathrm{\,GHz})\simeq 10^{40{\mbox{\scriptsize--}}41}\mathrm{\,erg\,s^{-1}}$ luminosities, and a preferentially radial radio-polarization. The narrow localization and properties of the excess identify these sources not as AGN, often invoked speculatively for excess X-ray sources at cluster outskirts, but rather as infalling gaseous clumps interacting with the virial shock, probably galactic halos and possibly outflow remnants. The local excess of such discrete, radio-to-$\gamma$-ray sources around an object can probe its virial shock also at high redshifts and sub-cluster scales.
@@ 1 Introduction
More recently, a $\sim 3\sigma$ excess of Chandra X-ray sources was reported [koulouridis2019high] in a fairly narrow, $2.52<\tau<$ 2.5 range of normalized radii around a small sample of five $z\sim 1$, mostly disturbed clusters. This signal is dominated by two clusters, each with about ten excess $2.52.0<\tau<$ 2.5 sources with respect to neighbouring bins. The excess sources show $10^{42.5}{\rm{\,erg\,s^{-1}}}\lesssim L_{0.5\mbox{ \scriptsize keV}}^{8.0\mbox{ \scriptsize keV}}\lesssim 10^{44}{\rm{\,erg\,s^{-1}}}$ luminosities, where energy or frequency subscripts (superscripts) denote low (high) band limits, henceforth. These sources were again tentatively attributed to enhanced AGN activity associated with excessive galaxy mergers or gas stripping.
@@
Therefore, a sharp spike in the radial profile of the density of sources would indicate a localized, sudden change in the environment. Interestingly, the structure formation, or virial, shock of a galaxy cluster, long anticipated to lie in the $32<\tau<$ 3 range according to cosmological simulations [e.g., KeshetEtAl03, SchaalSpringel15], provides a natural candidate for such a sharp environmental transition. Indeed, as an infalling object crosses the virial shock (VS), it should see the ambient pressure suddenly jump by orders of magnitude as the infalling surrounding plasma is violently slowed down and heated, accompanied by the generation of magnetic fields and shock-accelerated relativistic particles (cosmic rays, CRs) carrying a substantial fraction of the elevated thermal energy. While crossing a VS is unlikely to trigger an AGN, and the processes by which it could lead to the emergence of bright excess X-ray or radio sources may be a-priori unclear, the VS could explain at least the position and sharpness of the signal.
@@
The Supplementary material provides supporting and auxiliary information regarding the spectra of the excess X-ray sources (§A), limitations on source size estimates (§B), the evaluation of excess confidence levels (CLs) under Poisson statistics (§C), various convergence and sensitivity tests (§D), additional properties of the excess sources (§E), and the absence of a significant VS signal from 2RXS–NVSS pairs (§F) or identified AGN (§G). All (field+excess) sources found near the VS are listed in §H.
@@ 2.1 Cluster catalog and selection criteria
    Additional cuts are imposed on the Galactic latitude $|b|$, to avoid the bright Galactic plane, and on the declination, Dec. These cuts are specific to each source catalog, so are detailed below.
@@ 2.2 X-ray source catalog and cuts
Figure 1: Hammer-Aitoff all-sky projection of source density (cubehelix [Green11_Cubehelix] color map) and cluster sample (orange circles of radius $5{\theta_{500}}$), for X-ray (left panel; 2RXS) and radio (right panel, NVSS; note diminished exposure at low declinations, with no $\mbox{Dec}<-40^{\circ}$ sources) analyses. Also shown are the cuts of the Galactic plane (dotted yellow lines), the Fermi bubbles (dashed red curves), and Loop-I (dot-dashed cyan).
@@
Source fluxa cut $F^{*}$ $F_{X}<F_{X}^{*}\equiv 7\times 10^{-12}{\rm{\,erg\,s^{-1}\,cm^{-2}}}$ $F_{\nu}>F_{\nu}^{*}\equiv 100\mbox{ mJy}$
@@
Resultsb Mean field source density $f$ 5360 $5360\mbox{ sr}^{-1}$ 5760 $5760\mbox{ sr}^{-1}$
@@
  • a — Energy flux (specific energy flux) in X-rays (radio). The cuts are explained in the text.
    b — After cuts. Source numbers $\mathcal{N}$ are summed over all clusters. VS sources are listed in §H.
We define the X-ray energy flux $F_{X}\equiv F_{0.1\mbox{ \scriptsize keV}}^{2.4\mbox{ \scriptsize keV}}$ as the specific flux $F_{\epsilon}$ integrated over the ROSAT band. The catalog provides flux estimates based on three different emission models: a power-law spectrum, the optically-thin thermal plasma emission model Mekal [MeweEtAl86], and optically thick, black-body emission. In most or all relevant cases, the model fits are not sufficiently accurate to identify which emission model is most appropriate. Therefore, we henceforth adopt the power-law estimate of $F_{X}$, which is a simpler fit available for the largest number of sources; our results are not sensitive to this choice (see §D). We nevertheless consider below the fitted source temperatures, as well as the fitted photon power-law index $\Gamma_{X}$ defined by $F_{\epsilon}\propto\epsilon^{1+\Gamma_{X}}$.
Figure 2 (left panel) shows the energy flux histogram of X-ray sources in the analyzed sky region. The 2RXS catalog is dominated by foreground sources brighter than the stacked clusters themselves. To reduce foreground confusion, we impose an upper flux limit that removes sources whose luminosities would exceed the typical, $L_{X}\simeq 10^{45}{\rm{\,erg\,s^{-1}}}$ luminosity of the entire ICM even around the most distant ($z_{\rm max}\simeq 0.23$) clusters in our sample, thus retaining only $F_{X}<F_{X}^{*}\equiv 7\times 10^{-12}{\rm{\,erg\,s^{-1}\,cm^{-2}}}$ sources. Note that this $F_{X}^{*}$ cutoff is already quite close to the 2RXS $\sim 10^{-13}{\rm{\,erg\,s^{-1}\,cm^{-2}}}$ sensitivity threshold [2RXS_paper]. In §D, we show that our results are not sensitive to the precise choice of flux cut; for example, a high $F_{X}^{*}=10^{-10}{\rm{\,erg\,s^{-1}\,cm^{-2}}}$ changes the significance of excess virial sources by only $\sim 2\%$. In fact, a significant, $\sim 4\sigma$ signal persists even with no flux cut at all, despite the strong foreground.
@@ 2.3 Radio source catalog and cuts
The NVSS catalog covers the northern sky at declinations $\mbox{Dec}>-40^{\circ}$ (J2000). Figure 1 shows an apparent decline in $n_{\nu}$ with decreasing Dec even somewhat north of this declination. Therefore, we impose a corresponding cut on the declination of clusters used for radio stacking, nominally keeping only those with $\mbox{Dec}>-35^{\circ}$.
@@
A histogram of the specific flux of NVSS sources is shown in the right panel of figure 2. This catalog is strongly dominated by faint background sources (in contrast to the bright foreground source of 2RXS), due in part to star-forming galaxies, whose luminosity function peaks near $\nu L_{\nu}\sim 10^{39}{\rm{\,erg\,s^{-1}}}$ [e.g., radio_Lum]; for the most part, these sources are not expected to be associated with the low-redshift clusters in our sample. We therefore impose a lower specific-flux limit that removes sources with luminosities lower than this peak luminosity even around the nearest ($z_{\rm min}\simeq 0.017$) clusters in our sample, thus retaining only sources of $F_{\nu}\gtrsim F_{\nu}^{*}\equiv 100\mbox{ mJy}$. In §D, we show that changing the radio flux cutoff by $50\%$ has only a $\sim 15\%$ effect on the significance of our results, and our choice is conservative. Interestingly, the signal persists even if one avoids a flux cut entirely, thus retaining the large number of faint background sources, but expectedly diminishes to the $\sim 2\sigma$ level.
@@ 3 Stacking analysis method
$n(\tau,{\mathsf{c}})=\frac{\mathcal{N}(\tau,{\mathsf{c}})}{\Delta\omega(\tau,{\mathsf{c}})}\mbox{ ,}$ (3.1)
@@
$\Delta\omega(\tau,{\mathsf{c}})\equiv\frac{\Delta\Omega(\tau,{\mathsf{c}})}{\theta_{500,{\mathsf{c}}}^{2}}$ (3.2)
@@
$f(\tau,{\mathsf{c}})=\frac{\mathcal{F}(\tau,{\mathsf{c}})}{\Delta\omega(\tau,{\mathsf{c}})}\mbox{ ,}$ (3.3)
@@
$\delta(\tau,{\mathsf{c}})\equiv\frac{\mathcal{N}(\tau,{\mathsf{c}})-\mathcal{F}(\tau,{\mathsf{c}})}{\mathcal{F}(\tau,{\mathsf{c}})}=\frac{n(\tau,{\mathsf{c}})-f(\tau,{\mathsf{c}})}{f(\tau,{\mathsf{c}})}\mbox{ .}$ (3.4)
@@
$S\simeq\frac{\mathcal{N}-\mathcal{F}}{\sqrt{\mathcal{F}\,}}\mbox{ .}$ (3.5)
@@
$S_{\mbox{\tiny{joint}}}\simeq\frac{1}{\sqrt{N_{\mbox{\tiny{cat}}}}}\sum_{i}^{N_{\mbox{\tiny{cat}}}}S_{i},$ (3.6)
@@ 3.1 Source-weighted co-addition
$\mathcal{N}(\tau)\equiv\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\mathcal{N}(\tau,{\mathsf{c}})\mbox{ ,}$ (3.7)
@@
$\Omega(\tau)\equiv\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\Delta\Omega(\tau,{\mathsf{c}})\mbox{ ,}$ (3.8)
@@
$n_{\rm\mbox{\tiny{SW}}}(\tau)=\frac{\mathcal{N}(\tau)}{\bar{\omega}(\tau)}\mbox{ ,}$ (3.9)
@@
$\sigma[n_{\rm\mbox{\tiny{SW}}}(\tau)]\simeq\frac{\sqrt{\mathcal{N}(\tau)}}{\bar{\omega}(\tau)}\mbox{ .}$ (3.10)
Source-weighted estimates such as $n_{\rm\mbox{\tiny{SW}}}(\tau)$ are somewhat skewed toward extended clusters or clusters in strong-field regions, which may harbour a larger number of sources and so may dominate the numerator of Eq. (3.9). This bias is offset in part by the comparable extents of all clusters in our sample, and by avoiding regions with known contaminations.
The above approach is applied analogously to field sources, modelled with some distribution $\mathcal{F}(\tau,{\mathsf{c}})$ based on the field region. In analogy with Eqs. (3.7)–(3.9), the co-added number of field sources is
$\mathcal{F}(\tau)\equiv\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\mathcal{F}(\tau,{\mathsf{c}})\mbox{ ,}$ (3.11)
@@
$f_{\rm\mbox{\tiny{SW}}}(\tau)=\frac{\mathcal{F}(\tau)}{\bar{\omega}(\tau)}\mbox{ .}$ (3.12)
The significance of a local excess of sources with respect to the field-only null hypothesis can now be estimated as in Eq. (3.5),
$S_{\rm\mbox{\tiny{SW}}}(\tau)=\frac{\mathcal{N}(\tau)-\mathcal{F}(\tau)}{\sqrt{\mathcal{F}(\tau)}}\mbox{ ,}$ (3.13)
@@
$\delta_{\rm\mbox{\tiny{SW}}}(\tau)=\frac{n_{\rm\mbox{\tiny{SW}}}(\tau)}{f_{\rm\mbox{\tiny{SW}}}(\tau)}-1\mbox{ ,}$ (3.14)
@@ 3.2 Cluster-weighted co-addition
$S(\tau,{\mathsf{c}})\simeq\frac{\mathcal{N}(\tau,{\mathsf{c}})-\mathcal{F}(\tau,{\mathsf{c}})}{\sqrt{\mathcal{F}(\tau,{\mathsf{c}})}}\mbox{ .}$ (3.16)
@@
$S_{\rm\mbox{\tiny{CW}}}(\tau)=\frac{1}{\sqrt{{N}_{c}}}\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}S(\tau,{\mathsf{c}})\mbox{ .}$ (3.17)
@@
$n_{\rm\mbox{\tiny{CW}}}(\tau)=\frac{1}{{N}_{c}}\sum_{{\mathsf{c}}=1}^{{N}_{c}}n(\tau,{\mathsf{c}})\mbox{ ,}$ (3.18)
@@
$\sigma[n_{\rm\mbox{\tiny{CW}}}(\tau)]=\frac{1}{{N}_{c}}\sqrt{\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\frac{n(\tau,{\mathsf{c}})}{\Delta\omega(\tau,{\mathsf{c}})}}\mbox{ .}$ (3.19)
@@
$f_{\rm\mbox{\tiny{CW}}}(\tau)=\frac{1}{{N}_{c}}\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}f(\tau,{\mathsf{c}})\mbox{ .}$ (3.20)
@@
$\delta(\tau,{\mathsf{c}})=\frac{n(\tau,{\mathsf{c}})}{f(\tau,{\mathsf{c}})}-1\mbox{ ,}$ (3.21)
@@
$\delta_{\rm\mbox{\tiny{CW}}}(\tau)=\frac{1}{{N}_{c}}\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\delta\left(\tau,{\mathsf{c}}\right)\mbox{ ,}$ (3.22)
@@
$\sigma\left[\delta_{\rm\mbox{\tiny{CW}}}(\tau)\right]=\frac{1}{{N}_{c}}\sqrt{\sum\limits_{{\mathsf{c}}=1}^{{N}_{c}}\frac{n^{2}(\tau,{\mathsf{c}})}{f^{2}(\tau,{\mathsf{c}})\mathcal{N}(\tau,{\mathsf{c}})}}\mbox{ ,}$ (3.23)